# Overcoming Reward Model Noise in Instruction-Guided Reinforcement Learning

### v1

@@ 4 Approximation Error of Cosine Similarity
Cosine similarity is insufficient for evaluating the alignment between trajectories and instructions. In this section, we will discuss two fundamental issues associated with cosine similarity scores in RL contexts: state entanglement and composition insensitivity. The former issue, state entanglement, refers to the metric’s inability to recognize trajectories $\{\tau_{1},...,\tau_{n}\mid\tau_{i}\text{ induced by }\pi\in\Pi_{B}\}$ that, while being cosine similar to the target instruction in the embedding space, fail to reach the goal states in $S_{G}$. The latter issue refers to the metric’s tendency to reward trajectories that execute the sub-tasks (corresponding to individual instruction sentences $l_{k}$) in an incorrect order. This is problematic in RL tasks where the specific sequence of executing sub-tasks, as defined by the order of instructions in $L=\{l_{1},l_{2},\ldots,l_{n}\}$, is crucial for reaching the goal states.
4.1 The Issue of State Entanglement
@@
State entanglement refers to the issue where the cosine similarity metric erroneously pays more attention to lexical-level similarity while lacking comprehension of the underlying state transitions. Consequently, rewards are given to trajectory-instruction pairs that are cosine similar in embedding space but in fact result in distinct state transitions. For instance, consider the significant contrast between “take the red pill” and “take the blue pill”. Despite their lexical similarity, they lead to vastly different states. However, the cosine similarity metric may represent them as similar due to the shared action “take” and shared object “pill”, disregarding the critical difference in state outcomes. Understanding state transitions is crucial in sequential decision-making scenarios. Otherwise, rewards may be given to trajectories that lead to unintended states, resulting in the learning of unacceptable policies.
4.2 The Issue of Composition Insensitivity
@@
Composition insensitivity in cosine similarity metrics gives rise to two issues: (1) rewarding incomplete task execution – cosine similarity may incorrectly reward partial task completion, as even incomplete trajectories can receive high similarity score in the embedding space. For instance, in a task to “pick up a tool, then fix a machine,” the model might prematurely reward the agent for merely picking up the tool, neglecting the crucial repair action. This can lead to suboptimal learning, with the agent potentially getting stuck in easier, incomplete subtasks rather than fully accomplishing the task. (2) insensitivity to the ordering of execution – cosine similarity often fails to adequately penalize incorrect execution sequences. In a safety protocol requiring an agent to “turn off the machinery, then perform maintenance,” the metric might assign high rewards based merely on the presence of relevant actions, disregarding their order. While some large language models have become sensitive to the order of tokens (Su et al. 2024), the compact visual and sentence embeddings from multimodal models remains largely insensitive to sequential information (Pham et al. 2021). This limitation can lead to potentially dangerous outcomes in scenarios where execution sequence is critical, such as in safety protocols or complex manufacturing processes.
@@
Figure 1 illustrates various scenarios where high similarity scores are erroneously assigned to false positive cases. To empirically demonstrate the issue, Section 4.3 presents experiments on false positive rewards and their impact on agent learning in sparse reward environments.
@@
Figure 3: Learned VLM models differentiate between matched and not-matched pairs, but struggle with O.O.D. cases. They incorrectly assign high scores to manipulated pairs, which should be low as the trajectories in the manipulated pairs do not fulfill the instruction
@@ 4.3 Experiments on Reward Noise Impact
Our experiments test the following hypothesis: (H1) The two issues of state entanglement and composition insensitivity exist; (H2) false positive rewards are prevalent during training; (H3) similarity-based reward models lacking noise handling mechanisms underperform against strong intrinsic reward models in sparse reward environments; (H4) false negatives may not be as harmful as false positives.
@@
Setup. We evaluate these hypotheses through various challenging sparse-reward environments: (1) Crafter, an open-ended 2D Minecraft environment (Hafner 2021); (2) Montezuma’s Revenge, a classic hard adventure game in Atari (Bellemare et al. 2013); and (3) Minigrid ‘Go To Seq’, a hard task involving long-horizon navigation and object interactions (Chevalier-Boisvert et al. 2018). Two cosine similarity-based reward models were tested: (1) Pixl2R by Goyal et al. (2020), which uses only the current video frame to determine if the goal state specified in the instruction has been reached; and (2) ELLM-, a variant of ELLM by Du et al. (2023). Unlike ELLM, which queries instructions from LLMs in real-time, ELLM- directly uses preset expert instructions and compares them with the transition differences of the agent’s trajectory. The VLM backbones used are: (1) CLIP (Radford et al. 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al. 2022), pretrained by video-text pairs. To ensure high-quality finetuning data, we used internal information from the game engine to annotate expert trajectories from expert agents. To demonstrate how noisy reward signals hinder learning, we selected a strong intrinsic reward model DEIR (Zhang et al. 2021) for comparison. It provides auxiliary rewards based on observation novelty to encourage exploration. See Appendix A.3 for detailed implementation of the experiments.
@@
Evaluation Metric. We used the score metric adapted from the Crafter benchmark (Hafner 2021) for performance evaluation, as it better captures an agent’s consistent performance across multiple subtasks. This metric follows the intuitive ‘higher is better’ principle. Unlike maximum total rewards metric, which fail to capture consistent performance in sparse reward settings, the score metric provides a more reliable measure of learning progress.
@@
Reward Noise Issue. To investigate H1, we evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs. The state entanglement test involved reversing trajectories and negating instructions (i.e., “do not do $l_{k}$”). The composition insensitivity test examined concatenated pairs of trajectory-instruction data. For example, given $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we create a concatenated pair $(\tau_{1}+\tau_{2},l_{1}+l_{2})$. We then test two types of manipulations – (1) swapping the order within one modality: e.g., $(\tau_{2}+\tau_{1},l_{1}+l_{2})$; and (2) truncating one modality: e.g., $(\tau_{1},l_{1}+l_{2})$. Overall, we categorized the trajectory-instruction pairs for evaluation into three types: (1) matched pairs – these are positive examples where the trajectory correctly corresponds to the given instruction. (2) not-matched pairs – these are negative examples where the trajectory does not correspond to the given instruction. (3) manipulated pairs: derived from matched pairs by altering either the trajectory or the instruction. Ideally, manipulated pairs should receive low similarity scores as essentially the trajectory are not fulfilling the instruction. However, our results reveal that the reward model assigns high scores to these manipulated pairs, particularly in the “do not do $l_{k}$” case. This finding highlights the noise issue in cosine similarity-based reward models (see Figure 4). It’s worth noting that the poor performance in the negation case aligns with broader challenges in natural language processing. Recent studies (Hossain et al. 2022; Truong et al. 2023) have highlighted that negation is central to language understanding but is not properly captured by modern language models. This limitation extends to VLMs and significantly impacts their ability to provide accurate rewards in complex scenarios and instructions.
@@
Figure 5: The heatmap shows the cumulative rewards received at various locations, with larger circle sizes indicating higher rewards. The later figures shows the offsets between the state where rewards are given and the actual goal-reaching state. Agents are getting both issues of false positives and false negatives during training
@@
Prevalence of False Positives. To address H2, we analyzed reward distribution heatmap from VLM-based reward models during training. The heatmap revealed a concerning trend: RL agents engage in reward hacking, receiving rewards across vast areas of the environment rather than just at goal states. For instance, in Montezuma environment where the goal is to grab the key and escape the room, we observed that agents received rewards even for falling off cliffs, which undoubtedly contribute to the detrimental $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ specified in Theorem 3.1. For environments without fixed camera views, we calculated the step offset between the current rewarded state and the actual goal state. A positive offset indicates that a reward was given before reaching the goal state, signifying a false positive reward, while a negative offset suggests that the agent reached the goal state but the reward model failed to recognize it, indicating a false negative reward (see Figure 6). Interestingly, we observed a large amount of negative step offsets in Minigrid environments. We attribute this to Minigrid’s abstract shape-based visual representations, which fall outside the VLM’s pretraining distribution.
Impact on Learning. We trained agents using learned VLM reward models and compared their learning efficacy against intrinsic reward models. As shown in Table 1, our results confirmed H3: instruction-guided RL agents using learned VLM reward models without noise handling consistently underperform compared to DEIR, the intrinsic reward-based RL agent. To investigate the impact of false negatives versus false positives (H4), we designed an oracle Pixl2R model with two variants: a false negative model and a false positive model. The false negative model only rewards the agent for reaching subgoal states described in the instruction, with a probability of xthat some rewarding states in the map are removed. In contrast, the false positive model rewards the agent for reaching every subgoal, but also introduces a 0.1 one-off reward for certain locations, covering xof the map. The results indicate that false negatives may be less detrimental to agent performance than false positives (see Figure 4). This performance difference can be explained through our theoretical framework:
Table 1: Score metric across environments (equivalent to total rewards, higher is better). $\star$ denotes baseline intrinsic reward model. VLM reward models without noise handling underperformed. All models are based on PPO.
@@
PPO Pure $0.151$ $24.993$ $16.863$ $-28\%$
DEIR $\star$ Intrinsic $0.174$ $55.556$ $19.758$
Pixl2R VLM $0.142$ $12.422$ $9.409$ $\mathbf{-49\%}$
ELLM- VLM $0.150$ $19.406$ $10.826$ $-41\%$
Pixl2R + DEIR VLM + intr. $0.176$ $17.372$ $10.440$ $-38\%$
ELLM- + DEIR VLM + intr. $0.178$ $30.985$ $11.857$ $-27\%$
  1. 1.
    False negative models maintain $\mathbb{E}[G^{\pi\in\Pi_{G}}]-\mathbb{E}[G^{\pi\in\Pi_{B}}]$, ensuring steady gradient ascent towards maximizing $P(\pi\in\Pi_{G})$.
  2. 2.
    By decreasing $\mathbb{E}[G^{\pi\in\Pi_{B}}]$, false negative models minimize deviations from the target direction, leading to more stable learning.
In contrast, scenarios with high $\mathbb{E}[G^{\pi\in\Pi_{B}}]$ (i.e., the false positive case) significantly reduce the gradient ascent rate and also introduce deviations from the target direction. The findings thus challenged the prevailing assumption that false negatives are the primary issue in the degeneracy of instruction-guided RL. See Appendix A.4 for more experimental details.
5 Solution to the Reward Noise Issue
5.1 Preferring Binary Signal
@@
Our experiments have shown that the issue of false negatives may be less detrimental to learning than false positives. Therefore, we propose a reward function that provides one-time binary rewards only when the similarity between the agent’s current trajectory and the instruction exceeds a high confidence threshold. This method effectively reduces the likelihood of rewarding unacceptable trajectories. To achieve this, we introduce a thresholding mechanism using a calibration set of true positive trajectory-instruction pairs. This threshold, denoted as $\hat{q}$, is set to the empirical quantile of cosine similarity scores at the significance level $1-\alpha$. Pairs whose similarity scores fall below this threshold $\hat{q}$ receive no reward. Conversely, pairs exceeding $\hat{q}$ receive a one-time $+1$ reward, i.e., $r^{v}_{\textsc{Bi}}(\tau,l_{k})=\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}$. This approach statistically guarantees a high probability (at least $1-\alpha$) that true positive pairs are recognized while minimizing frequency of false positives errors (Sadinle et al. 2019). See Appendix A.5 for detailed threshold calculation.
@@ 5.2 Mutual Information Maximization
$=\mathbb{E}_{\tau\sim\pi_{\theta},l_{k}\sim L}[\log p(l_{k}\mid\tau)-\log p(l_{k})]$ (1)
where $p(l_{k}\mid\tau)$ comes from the similarity score provided by VLMs, and $p(l_{k})$ is the frequency of the instruction $l_{k}$ being fulfilled by the agent’s trajectory. Therefore, the second term in the equation serves as a regularization term that downplays the significance of the reward signal when it is too frequent. For instance, if a VLM frequently detects that the agent’s actions are fulfilling the “climbing the ladder” instruction, even when the agent is performing unrelated tasks, the reward signal for this instruction will be downplayed. $p(l_{k})$ is calculated as follows:
$p(l_{k})=\mathbb{E}_{\tau\sim\pi_{\theta_{-1}}}\left[\frac{\sum_{t=1}^{T_{\tau}}\mathbf{1}_{\{p(l_{k}\mid\tau_{t})\geq\hat{q}\}}}{T_{\tau}}\right]$ (2)
@@
It’s important to note that the BiMI approach primarily mitigates false positives (FP) rather than false negatives (FN). By implementing a high confidence threshold and a binary reward signal, we effectively reduce the likelihood of rewarding incorrect trajectories, thus addressing the FP issue. While this conservative approach may potentially increase FNs, as some correct trajectories might fall below the threshold, our experiments indicate that this trade-off is beneficial. We found that false positives are more detrimental to learning in instruction-guided RL tasks, justifying our preference for a method that prioritizes FP reduction.
In the following sections, we will demonstrate how BiMI significantly improves the effectiveness of VLM-based reward models in instruction-guided RL tasks.

### v4

@@ 5 False Positives From Cosine Similarity
This section identifies and discusses two fundamental issues with cosine similarity scores in sequential decision-making contexts: state entanglement and composition insensitivity. The former issue, state entanglement, refers to the metric’s inability to recognize trajectories that, while being cosine similar to the target instruction in the embedding space, fail to reach the goal states in $S_{G}$. The latter issue refers to the metric’s tendency to reward trajectories even when the temporal relationships between sub-tasks are not satisfied.
@@
The Issue of State Entanglement State entanglement refers to the issue where the cosine similarity metric erroneously pays more attention to lexical-level similarity while lacking comprehension of the underlying state transitions. Consequently, rewards are given to trajectory-instruction pairs that are cosine similar in embedding space but in fact result in distinct state transitions. For instance, consider the significant contrast between “take the red pill” and “take the blue pill”. Despite their lexical similarity, they lead to vastly different states. However, the cosine similarity metric may represent them as similar due to the shared words, disregarding the critical difference in state outcomes. Understanding state transitions is crucial in sequential decision-making scenarios. Otherwise, rewards may be given to trajectories that lead to unintended states, potentially prolonging the path to the goal state by necessitating corrective actions or re-attempts.
@@
The Issue of Composition Insensitivity Composition insensitivity in cosine similarity metrics gives rise to two issues: (1) rewarding incomplete task execution – cosine similarity may incorrectly reward partial task completion, as even incomplete trajectories can receive high similarity score in the embedding space. For instance, in a task to “pick up a tool, then fix a machine,” the model might prematurely reward the agent for merely picking up the tool, neglecting the crucial repair action. We also observed this phenomenon particularly in the Montezuma environment, where RL agents tend to hack the reward system by focusing on the easiest actions that yield rewards (e.g., moving towards a direction) rather than executing more complex, timely actions. Eventually, this leads to an overestimation of the agent’s progress towards the ultimate goal. (2) insensitivity to the ordering of execution – cosine similarity often fails to adequately penalize incorrect execution sequences. In a safety protocol requiring an agent to “turn off the machinery, then perform maintenance,” the metric might assign high rewards based merely on the presence of relevant actions, disregarding their order. In contrast to some advancements in language models, compact visual and sentence embeddings from multimodal VLMs remain largely insensitive to sequential information (Pham et al., 2021). When the task is order-sensitive, executing actions in the wrong sequence prolongs the path towards the goal state, as agents need to re-attempt the correct order.
@@
Figure 1 illustrates various scenarios where false positive rewards are erroneously assigned. To empirically demonstrate the issue, Section 5.1 presents experiments on these issues and their impact on agent learning in sparse reward environments.
@@
Figure 3: Learned VLM models performed badly with O.O.D. examples. They incorrectly assign high scores to manipulated pairs, which should be low as the trajectories in the manipulated pairs fail the instruction.
@@ 5.1 Experiments on Reward Noise Impact
Our experiments test the following hypothesis: (H1) The two issues of state entanglement and composition insensitivity exist; (H2) false positive rewards are prevalent during training; (H3) VLM reward models lacking noise handling mechanisms underperform against intrinsic reward models in sparse reward environments; (H4) false negatives may not be as harmful as false positives.
@@
Setup. We evaluate these hypotheses through various challenging sparse-reward environments: (1) Crafter, an open-ended 2D Minecraft (Hafner, 2021); (2) Montezuma, a classic hard adventure game in Atari (Bellemare et al., 2013); and (3) Minigrid ‘Go To Seq’, a hard task involving long-horizon navigation and object interactions (Chevalier-Boisvert et al., 2018). A Markovian and a Non-Markovian reward model were tested: (1) Pixl2R by Goyal et al. (2020), which uses only the current video frame to determine if the goal state specified in the instruction has been reached; and (2) ELLM-, a variant of ELLM by Du et al. (2023). Unlike ELLM, which queries instructions from LLMs in real-time, ELLM- directly uses preset expert instructions and compares them with the transition differences of the agent’s trajectory. The VLM backbones used are: (1) CLIP (Radford et al., 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al., 2022), pretrained by video-text pairs. To ensure high-quality finetuning data, we used internal information from the game engine to annotate expert trajectories from expert agents. To demonstrate how noisy reward signals hinder learning, we selected a strong intrinsic reward model DEIR (Zhang et al., 2021) for comparison. It provides auxiliary rewards based on observation novelty to encourage exploration. See Appendix A.4 for detailed implementation of the experiments.
@@
Evaluation Metric. We adopted the score metric from the Crafter benchmark (Hafner, 2021) for performance evaluation, as it effectively measures consistent performance across multiple subtasks in sparse reward environments. Unlike the maximum total rewards metric, which does not adequately reflect consistent performance, the score metric offers a more reliable indicator of learning progress. See Appendix A.5 for its formal definition.
@@
Reward Noise Issue. To investigate H1, we evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs. The state entanglement test involved reversing trajectories and negating instructions (i.e., “do not do $l_{k}$”). The composition insensitivity test examined concatenated pairs of trajectory-instruction data. Given $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we create a concatenated pair $(\tau_{1}+\tau_{2},l_{1}+l_{2})$. We then test two types of manipulations – (1) swapping the order within one modality: e.g., $(\tau_{2}+\tau_{1},l_{1}+l_{2})$; and (2) truncating one modality: e.g., $(\tau_{1},l_{1}+l_{2})$. Overall, three types are evaluated: (1) matched pairs; (2) not-matched pairs and (3) manipulated pairs which are derived from matched pairs by polluting either the trajectory or the instruction. Our results reveal a critical flaw in the reward model: despite manipulated pairs that fundamentally fail to fulfill the instruction, the model paradoxically assigns high similarity scores (see Figure 4 for overall results and also Appendix A.6 for individual environments). It’s worth noting that the poor performance in the negation case aligns with broader challenges in natural language processing. Recent studies (Hossain et al., 2022; Truong et al., 2023) have highlighted that negation is central to language understanding but is not properly captured by modern language models. This limitation extends to VLMs and directly leads to false positive rewards.
@@
Figure 5: The heatmap shows rewards received at various locations, with larger circle sizes indicating higher rewards. The later figures shows the offsets between the state where rewards are given and the actual goal-reaching state. Agents are getting both issues of false positives and false negatives
@@
Table 1: Score metric across environments (equivalent to total rewards, higher is better). $\star$ denotes baseline intrinsic reward model. VLM reward models without noise handling underperformed. All are based on PPO.
@@
PPO Pure $0.151$ $24.9$ $16.8$ $-28\%$
DEIR $\star$ Intrinsic $0.174$ $55.5$ $19.7$
Pixl2R VLM $0.142$ $12.4$ $9.40$ $\mathbf{-49\%}$
ELLM- VLM $0.150$ $19.4$ $10.8$ $-41\%$
Pixl2R + DEIR VLM + intr. $0.176$ $17.3$ $10.4$ $-38\%$
ELLM- + DEIR VLM + intr. $0.178$ $30.9$ $11.8$ $-27\%$
Prevalence of False Positives. To address H2, we analyzed reward distribution heatmap from VLM-based reward models during training. The heatmap revealed a concerning trend: RL agents engage in reward hacking, receiving rewards across vast areas of the environment rather than just at goal states. For instance, in Montezuma where the goal is to grab the key and escape the room, we observed that agents received rewards even for falling off cliffs, which undoubtedly contribute to false positive rewards. For environments without fixed camera views, we calculated the step offset between the current rewarded state and the actual goal state. A positive offset indicates a false positive reward, as the reward was given before reaching the goal. Conversely, a negative offset indicates a false negative, where the agent reached the goal but the reward model failed to acknowledge it (see Figure 6). Interestingly, besides positive offsets, we observed a large amount of negative offsets in Minigrid environments. We attribute this to Minigrid’s abstract shape-based visual representations, which fall outside the VLM’s training distribution.
Impact on Learning. We trained agents using learned VLM reward models and compared their learning efficacy against intrinsic reward models. As shown in Table 1, our results confirmed H3: instruction-following RL agents using learned VLM reward models without noise handling consistently underperform compared to DEIR, the intrinsic reward-based RL agent. To investigate the impact of false negatives versus false positives (H4), we designed an oracle Pixl2R model with two variants: a false negative model and a false positive model. The false negative model only rewards the agent for reaching subgoal states described in the instruction, with a probability of xthat some rewarding states in the map are removed. In contrast, the false positive model rewards the agent for reaching every subgoal, but also introduces a small (0.1) one-off reward for certain locations, covering xof the map. The results indicate that false negatives were less detrimental to agent performance than false positives (see Figure 4). This performance difference can be explained through our theoretical analysis in Appendix A.2.2.
6 Addressing the False Positive Issue
Our proposed solution is not specific to cosine similarity’s approximation error, which serves as a case study. Rather, it’s a broad strategy for mitigating false positive rewards across various reward models, including noisy VLM reward sources.
6.1 Binary Signal and Conformal Prediction
Our experiments have demonstrated that false positives are more detrimental to learning than false negatives. Based on these findings, we propose a reward function that issues a one-time binary reward only when the similarity between the agent’s current trajectory and the instruction exceeds a high confidence threshold. This approach contrasts with previous methods, which provide continuous rewards whenever the reward score exceeds a predefined threshold, and continue to do so until reaching a maximum cap. Our method, however, delivers this reward only once. This approach minimizes the likelihood of accumulating false positive rewards while maintaining adherence to Assumption 4.3.
@@ 6.1 Binary Signal and Conformal Prediction
To achieve this, we introduce a thresholding mechanism using a calibration set of true positive trajectory-instruction pairs. This threshold, denoted as $\hat{q}$, is set to the empirical quantile of cosine similarity scores at the significance level $1-\alpha$. Pairs whose similarity scores fall below this threshold $\hat{q}$ receive no reward. Conversely, pairs exceeding $\hat{q}$ receive a one-time $+1$ reward, i.e., $r^{v}_{\textsc{Bi}}(\tau,l_{k})=\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}$. This thresholding approach leverages statistical properties studied by Sadinle et al. (2019), which ensures a high probability (at least $1-\alpha$) that true positive pairs are recognized while minimizing frequency of false positives errors. See Appendix A.4.2 for the pseudo-code of the instruction-following RL algorithm and Appendix A.8 for detailed threshold calculation.
@@ 6.2 Mutual Information Maximization
$=\mathbb{E}_{\tau\sim\pi_{\theta},l_{k}\sim L}[\log p(l_{k}\mid\tau)-\log p(l_{k})]$
where $\tau=\langle s_{t-W},a_{t-W},\ldots,s_{t}\rangle$ is the agent’s trajectory up to current time step $t$, and $W$ is the memory size of the agent for its past trajectory. $p(l_{k}\mid\tau)$ comes from the similarity score provided by VLMs, referring to the likelihood of the instruction $l_{k}$ being fulfilled by a trajectory $\tau$. $p(l_{k})$ is overall likelihood of encountering the instruction $l_{k}$ in the learning environment. Therefore, the second term in the equation serves as a regularization term that downplays the significance of the reward signal when it is too frequent. For instance, if a VLM frequently detects that the agent’s actions are fulfilling the “climbing the ladder” instruction, even when the agent is performing unrelated tasks, any reward signal from this instruction will be downplayed. $p(l_{k})$ is calculated as follows:
$p(l_{k})=\mathbb{E}_{\tau\sim\pi_{\theta_{-1}}}\left[\left(\sum\nolimits_{t=1}^{T_{\tau}}\mathbf{1}_{\{p(l_{k}\mid\tau_{t})\geq\hat{q}\}}\right)/T_{\tau}\right]$
@@
Methods Montezuma % vs. $\star$ Minigrid % vs. $\star$ Crafter % vs. $\star$
Pixl2R $\star$ $0.142\pm 0.003$ $12.4\pm 2.43$ $9.40\pm 0.022$
Pixl2R + Bi $0.137\pm 0.009$ $-3.5\%$ $31.2\pm 2.04$ $+151\%$ $10.7\pm 0.784$ $+14\%$
Pixl2R + BiMI $0.162\pm 0.022$ $+14\%$ $37.5\pm 7.83$ $+199\%$ $7.95\pm 0.351$ $-15\%$
Pixl2R + DEIR $0.176\pm 0.009$ $+23\%$ $17.3\pm 0.51$ $+39\%$ $10.4\pm 1.015$ $+10\%$
Pixl2R + BiMI + DEIR $0.267\pm 0.016$ $+88\%$ $57.7\pm 2.15$ $+364\%$ $11.0\pm 0.190$ $+17\%$
ELLM- $\star$ $0.150\pm 0.004$ $19.4\pm 10.06$ $10.8\pm 1.017$
ELLM- + Bi $0.151\pm 0.016$ $+0.6\%$ $29.7\pm 1.29$ $+53\%$ $11.1\pm 0.601$ $+3.2\%$
ELLM- + BiMI $0.156\pm 0.014$ $+4.0\%$ $33.6\pm 3.99$ $+74\%$ $9.42\pm 0.267$ $-12\%$
ELLM + DEIR $0.178\pm 0.029$ $+20\%$ $30.9\pm 3.50$ $+59\%$ $11.8\pm 1.152$ $+9.5\%$
ELLM- + BiMI + DEIR $0.279\pm 0.078$ $+86\%$ $56.2\pm 6.19$ $+190\%$ $13.1\pm 0.393$ $+22\%$
@@
It’s important to note that the BiMI approach primarily mitigates false positives (FP) rather than false negatives (FN). Both Bi and MI aim to reduce the likelihood of rewarding unintended trajectories, thus addressing the FP issue. While this conservative approach may increase false negatives by pruning some correct trajectories, this trade-off is beneficial, as demonstrated by both our empirical and theoretical results.